OpenStudy (experimentx):

Show that the determinant and trace of Matrix remains invariant under similarity transformation

OpenStudy (anonymous):

\[ det( A^{-1} B A) =det(A^{-1}) det(B) det(A)=\frac 1 {det(A)}det(B) det (A) =det(B) \] Can you do the trace?

OpenStudy (anonymous):

Use that trace(MN)=trace(NM)

OpenStudy (anonymous):

\[ trace(A^{-1} B A)=trace( A^{-1} A B)=trace(B) \]

OpenStudy (experimentx):

thanks ... prof